OpenStudy (anonymous):

solve by completing the square x^2 - 4x + 2 = 0

OpenStudy (rogue):

\[x^2 - 4x + 2 = 0\]\[x^2 - 4x = -2\]\[x^2 - 4x + \frac {(-4)^{2}}{4}= -2 + \frac {(-4)^{2}}{4}\]\[x^2 - 4x + 4 = -2 + 4\]\[(x - 2)^2 = 2\]\[x - 2 = \pm \sqrt {2}\]\[x = 2 \pm \sqrt 2\]